# Uncertainty-Aware Collaborative System of Large and Small Models for Multimodal Sentiment Analysis

### v1

@@ Appendix A Appendix
Appendix B B  Uncertainty-aware Baseline Model (UBM) Architecture
This section provides detailed information about the architecture of UBM, which serves as the efficient filtering component in our U-ACS framework.
Overall UBM Architecture
The UBM is built upon a multimodal sentiment analysis framework with key modifications to support uncertainty estimation and improve multimodal fusion capabilities. The model consists of three main components: modality-specific encoders, multimodal fusion mechanisms, and dual prediction heads (regression and classification).
Feature Extraction
(1) Visual Modality: For the CMU-MOSI and CMU-MOSEI datasets, following our baseline Self-MM (Yu et al. 2021), Facet111iMotions 2017. https://imotions.com is utilized to extract a collection of visual features, which include facial action units, facial landmarks, head pose, gaze tracking, and histogram of oriented gradients (HOG) features, among others. These visual features are extracted from each utterance at a rate of 30Hz, resulting in a temporal sequence of facial gestures. For the CH-SIMS dataset, MTCNN face detection algorithm (Zhang et al. 2016) is first applied to obtain aligned faces. Subsequently, the MultiComp OpenFace2.0 toolkit (Baltrusaitis et al. 2018) is employed to extract a set of 68 facial landmarks, 17 facial action units, head pose, head orientation, and eye gaze, among other features.
(2) Acoustic Modality: For the CMU-MOSI and CMU-MOSEI datasets, audio analysis framework COVAREP (Degottex et al. 2014) is employed to extract acoustic features. These features include 12 Mel-frequency cepstral coefficients, pitch tracking, speech polarity, glottal closure instants, spectral envelope, peak slope parameters, and maxima dispersion quotients, among others. The acoustic features are extracted from the full audio clip of each utterance at a sampling rate of 100Hz, forming a sequence that captures variations in the tone of voice throughout the utterance. For the CH-SIMS dataset, LibROSA (McFee et al. 2015) speech toolkit, configured with default parameters, is utilized to extract acoustic features at a sampling rate of 22,050Hz. The extracted features consist of 33-dimensional frame-level acoustic features, which include 1-dimensional logarithmic fundamental frequency (log F0), 20-dimensional Mel-frequency cepstral coefficients (MFCCs), and 12-dimensional Constant-Q chromatogram (CQT).
Modality-specific Encoders
Text Encoder: For textual input processing, we employ BERT as the backbone encoder. The model supports both English (BERT-base-uncased) and Chinese (BERT-Chinese) variants depending on the dataset language. Text sequences are first tokenized using the BERT tokenizer, then fed into the pre-trained BERT model:
$\begin{split}\bm{H}_{t}&=\text{BERT}(\bm{U}_{t})\in\mathbb{R}^{T\times d_{bert}},\\ \bm{x}_{t}&=\bm{H}_{t}[0,:]\in\mathbb{R}^{d_{bert}}\end{split}$ (B.1)
where $\bm{U}_{t}$ represents the tokenized text input with shape $(B,3,T)$ containing input_ids, attention_mask, and token_type_ids, $\bm{H}_{t}$ is the contextualized hidden states from BERT, and $\bm{x}_{t}$ is the final text representation extracted from the [CLS] token. The text output dimension $d_{bert}=768$ is fixed across all datasets.
Audio Encoder: For acoustic feature processing, for the sake of efficiency, follow Self-MM (Yu et al. 2021), we utilize an LSTM-based subnet to capture temporal dependencies, which has few parameters:
$\begin{split}\text{packed\_audio}&=\text{pack\_padded\_sequence}(\bm{U}_{a},\text{lengths}),\\ \_,(\bm{h}_{final},\_)&=\text{LSTM}(\text{packed\_audio};\theta_{a}),\\ \bm{x}_{a}&=\text{Linear}(\text{Dropout}(\bm{h}_{final}))\end{split}$ (B.2)
where $\bm{U}_{a}\in\mathbb{R}^{T\times f_{a}}$ represents the input audio features, and $\bm{x}_{a}\in\mathbb{R}^{d_{a\_out}}$ is the final audio representation with $d_{a\_out}=16$.
$\begin{split}\text{packed\_video}&=\text{pack\_padded\_sequence}(\bm{U}_{v},\text{lengths}),\\ \_,(\bm{h}_{final},\_)&=\text{LSTM}(\text{packed\_video};\theta_{v}),\\ \bm{x}_{v}&=\text{Linear}(\text{Dropout}(\bm{h}_{final}))\end{split}$ (B.3)
where $\bm{U}_{v}\in\mathbb{R}^{T\times f_{v}}$ represents the input visual features, and $\bm{x}_{v}\in\mathbb{R}^{d_{v\_out}}$ is the final visual representation with $d_{v\_out}=32$.
Multimodal Feature Fusion
The extracted unimodal representations are concatenated and processed through a fusion network:
$\begin{split}\bm{x}_{concat}&=\bm{x}_{t}\|\bm{x}_{a}\|\bm{x}_{v},\\ \bm{h}_{m}&=\text{Dropout}(\text{ReLU}(W_{1}\bm{x}_{concat}+b_{1}))\end{split}$ (B.5)
where the concatenated dimension is 816 (768+16+32), and $\bm{h}_{m}\in\mathbb{R}^{128}$ is the fused representation.
Dual Prediction Heads
Regression Predictor: Predicts continuous sentiment scores:
$\hat{y}_{s}=W_{r}^{(2)}\text{ReLU}(W_{r}^{(1)}\bm{h}_{m}+b_{r}^{(1)})+b_{r}^{(2)}$ (B.6)
Classification Predictor: Performs 3-class sentiment classification:
$\begin{split}\bm{z}_{cls}&=W_{c}^{(2)}\text{ReLU}(W_{c}^{(1)}\bm{h}_{m}+b_{c}^{(1)})+b_{c}^{(2)},\\ \bm{p}_{s}&=\text{Softmax}(\bm{z}_{cls})\end{split}$ (B.7)
Training Objectives
The model is trained with a combined loss function:
$\mathcal{L}_{total}=\text{MSE}(\hat{y}_{s},y)+\text{CE}(\bm{p}_{s},y_{cls})$ (B.8)
where $y_{cls}$ is the discretized sentiment label defined as:
$y_{cls}=\begin{cases}1\;(\text{Negative}),&y<0\\ 0\;(\text{Neutral}),&y=0\\ 2\;(\text{Positive}),&y>0\end{cases}$ (B.9)
Uncertainty Estimation
During inference, uncertainty is computed as the entropy of the classification probability distribution:
$u_{s}=-\sum_{i=1}^{3}p_{s,i}\cdot\log(p_{s,i}+\epsilon)$ (B.10)
where $\epsilon=10^{-12}$ prevents numerical instability.
Dataset-specific Configurations
The input feature dimensions and LSTM hidden sizes vary across datasets:
  • MOSI: Feature dims: (768, 5, 20), LSTM hidden: audio=32, video=64
  • MOSEI: Feature dims: (768, 74, 35), LSTM hidden: audio=32, video=32
  • SIMS: Feature dims: (768, 33, 709), LSTM hidden: audio=16, video=64
Optimization Strategy
@@ Appendix B B  Uncertainty-aware Baseline Model (UBM) Architecture
UBM is trained using the Pytorch Framework. The model employs component-specific learning rates and weight decay values that vary across datasets. Table B.1 presents the detailed optimization parameters for each dataset.
@@
Weight Decay $1\times 10^{-3}$ $0.0$ $1\times 10^{-2}$
@@
Weight Decay $1\times 10^{-3}$ $0.0$ $1\times 10^{-2}$
@@
Table B.1: Component-specific optimization configurations for each dataset.
All components use the Adam optimizer with component-specific parameter groups to enable fine-grained control over the training dynamics of different model parts.
Appendix C C  Experimental Details and Additional Results
In this section, we provide additional experimental details and additional results, including the a more comprehensive comparison between the proposed three uncertainty quantification methods, the ablation studies on CMU-MOSIE datasets. These results are consistent with the results on the main text, further demonstrating the effectiveness of the proposed method.
Dataset Description Acc7 Acc2 MAE
MOSEI Prediction-Truth Difference 54.4 84.7 0.520
Ensemble Variance 55.2 86.4 0.516
Ours 55.4 86.4 0.506
Table B.2: Ablation experiments on CMU-MOSEI. Best results are in bold, second-best are underlined.
@@ Appendix C C  Experimental Details and Additional Results
PTD 45.5 85.7 85.6 0.721 0.791
EV 43.6 84.5 84.4 0.743 0.784
CBE 45.9 85.8 85.7 0.732 0.786
Table B.3: Comparison of uncertainty quantification methods on CMU-MOSI dataset. PTD: Prediction-Truth Difference, EV: Ensemble Variance, CBE: Classification-based Entropy (Ours).
Appendix D D  Prompt Design for MLLM Integration
Dataset-specific Prompt Templates
CH-SIMS v1 Dataset
Sentiment scores range from -1 to +1, where -1 is highly negative, +1 is highly positive, and 0 is neutral. The speaker said ‘{text}’. What is the sentiment score of the person in the video based on visual, audio and text? Directly answer the sentiment score. Now, the model is required to re-predict the label of the incoming data, and at the same time, provide the results of uncertainty and predicted labels obtained by the BERT-based multimodal sentiment prediction model and MLLM: {data}.
CMU-MOSI and CMU-MOSEI Datasets
Sentiment scores range from -3 to +3, where -3 is highly negative, +3 is highly positive, and 0 is neutral. The speaker said ‘{text}’. What is the sentiment score of the person in the video based on visual, audio and text? Directly answer the sentiment score. Now, the model is required to re-predict the label of the incoming data, and at the same time, provide the results of uncertainty and predicted labels obtained by the BERT-based multimodal sentiment prediction model and MLLM: {data}.
Template Variables
The prompt templates use the following variable substitutions:
  • {text}: Transcribed text content from the video segment
  • {data}: Structured data containing predictions and uncertainties from both UBM and MLLM models
Design Principles
Our prompts follow a two-stage approach: (1) direct multimodal sentiment prediction, and (2) collaborative integration of both small and large model predictions with uncertainty estimates. Different sentiment scales are used: CH-SIMS uses a 7-point scale (-3 to +3), while MOSI/MOSEI use a continuous scale (-1 to +1).

### v2

@@ IV-B3 Efficiency Metrics
IV-C Baseline and Compared Models
This section describes the baseline models used in our experiments, including MISA, HumanOmni, VideoLLaMA2 and so on.
IV-C1 MISA [13]
Learns modality-invariant (shared) and modality-specific (unique) representations from multimodal data.
IV-C2 Low-rank Multimodal Fusion (LMF) [23]
Employs low-rank tensors for efficient multimodal fusion, significantly reducing computational complexity compared to traditional tensor-based methods.
IV-C3 Multimodal Transformer (MulT) [41]
A Multimodal Transformer designed to handle unaligned multimodal sequences through the use of directional pairwise cross-modal transformers.
IV-C4 Multimodal Adaptation Gate-BERT (MAG-BERT) [36]
Integrates non-verbal modalities (audio and visual) into pre-trained language models like BERT using a multimodal attention gating mechanism.
IV-C5 Self-Supervised Multi-task Multimodal sentiment analysis network (Self-MM) [59]
A self-supervised framework that automatically generates unimodal labels to guide a multi-task learning process, avoiding the need for manual unimodal annotation.
IV-C6 Contrastive FEature DEcomposition framework (ConFEDE) [55]
Decomposes features from each modality into similarity and dissimilarity components using a unified contrastive learning framework.
IV-C7 Disentangled-Language-Focused (DLF) multimodal representation learning framework [44]
A framework that prioritizes the language modality by disentangling shared and specific features and using a Language-Focused Attractor to leverage complementary information.
IV-C8 Tensor Fusion Network (TFN) [60]
A pioneering model that uses tensor fusion via outer products to capture high-order, complex inter-modal dynamics.
IV-C9 Knowledge-Guided Dynamic Modality Attention Fusion Framework (KuDA) [9]
Dynamically selects the dominant modality for a given sample using sentiment knowledge, rather than statically assuming text is always the most important.
IV-C10 VideoLLaMA2 [5]
A powerful MLLM with strong spatial-temporal and audio understanding capabilities. It is used as a standalone baseline to benchmark the efficiency and performance of our U-ACS framework.
IV-C11 HumanOmni [64]
A 7-billion-parameter MLLM that serves as the expert oracle within our U-ACS framework. It is LoRA-adapted to process high-uncertainty samples that are escalated from the baseline model.
IV-D Experimental Details
We develop our approach using PyTorch on NVIDIA GeForce RTX 4090 with CUDA 13.0 and torch 2.2.1.
@@ IV-D Experimental Details
UBM is trained using the PyTorch Framework. The model employs component-specific learning rates and weight decay values that vary across datasets. Table I presents the detailed optimization parameters for each dataset.
For HumanOmni and VideoLLaMA2, we configure the training epochs, learning rate, LoRA Rank, and LoRA $\alpha$ to 1, 2e-5, 128, and 256 respectively.
TABLE I: Component-specific optimization configurations.
@@
Weight Decay $1\times 10^{-3}$ 0 $1\times 10^{-2}$
@@
Weight Decay $1\times 10^{-3}$ 0 $1\times 10^{-2}$
@@
IV-E Result on Multimodal Sentiment Analysis
To assess the performance of U-ACS, we utilized three well-recognized MSA datasets. HumanOmni is employed as the core MLLM in U-ACS. As presented in Table II, U-ACS demonstrates superior performance, achieving state-of-the-art results in some evaluation metrics while maintaining computational efficiency. Compared to HumanOmni, U-ACS achieves 0.5 points improvement in Acc2 and F1 score on MOSI, and achieves 0.3 points improvement in Acc2, 0.4 points improvement in F1 score on MOSEI. Our approach achieves superior performance over conventional multimodal methods on all metrics. Table III presents the robustness validation results of U-ACS in cross-lingual scenarios. On SIMS, the model maintains leading performance across most metrics, including Acc3, Acc2 and F1-score. Furthermore, U-ACS significantly outperforms all traditional multimodal methods, demonstrating the effectiveness of uncertainty-guided collaboration across different languages and data characteristics.
It is noteworthy that while U-ACS consistently improves binary classification metrics, the standalone MLLM occasionally retain an edge in regression-focused metrics like MAE and Corr. This outcome is not an unintended side effect but a direct consequence of our framework’s dual design philosophy: to maximize both classification accuracy and computational efficiency. Our core uncertainty estimation and sample routing mechanisms are designed precisely to serve these two goals, which also explains the trade-off in regression performance. Specifically, the UBM estimates uncertainty by converting the regression task into a simpler three-class classification problem and calculating prediction entropy. Furthermore, the routing thresholds are determined by evaluating whether the predicted polarity aligns with the ground truth, a process that is both fast and directly tied to the primary goal of Acc2. The entire cascade is therefore architected to quickly filter the majority of ‘easy’ samples using a cheap model, drastically reducing the average inference time and reserving the expensive MLLM only for challenging cases. This design creates a deliberate trade-off. The standalone MLLM, which processes every sample, is singularly focused on minimizing the error of fine-grained sentiment scores across the board, thus naturally excelling at metrics like MAE and Corr that measure the precision of continuous values. Our U-ACS, in contrast, prioritizes rapidly and correctly identifying the sentiment polarity for the bulk of the data. It will accept a less precise regression value from the small model if the polarity confidence is high, because doing so is far more efficient. Consequently, we achieve superior performance in Acc2 and F1-scores and a significant reduction in runtime, at the minor, and acceptable, cost of precision in sentiment intensity prediction for some samples.
TABLE II: Performance comparison on MOSI and MOSEI datasets. Best results are in bold, second-best are underlined. means the results are from [9].
Method MOSI MOSEI
@@ IV-E Result on Multimodal Sentiment Analysis
MISA [13] 42.3 83.4 83.6 0.783 0.761 52.2 85.5 85.3 0.555 0.756
LMF [23] 33.8 79.2 79.2 0.950 0.651 51.6 83.5 83.4 0.575 0.716
MulT [41] 36.9 79.7 79.6 0.879 0.702 52.8 84.6 84.5 0.559 0.733
MAG-BERT [36] 43.6 84.4 84.6 0.727 0.781 52.7 84.8 84.7 0.543 0.755
Self-MM [59] 45.8 84.9 84.8 0.731 0.785 53.0 85.2 85.2 0.540 0.763
ConFEDE [55] 42.3 85.5 85.5 0.742 0.784 54.9 85.8 85.8 0.522 0.780
DLF [44] 47.1 85.1 85.0 0.731 0.781 53.9 85.4 85.3 0.536 0.764
HumanOmni [64] 52.8 91.3 91.3 0.549 0.881 58.6 86.1 85.4 0.483 0.807
U-ACS 51.2 91.8 91.8 0.585 0.868 55.4 86.4 85.8 0.506 0.784
TABLE III: Performance comparison on SIMS dataset. Best results are in bold, second-best are underlined. means the results are from [9].
Method SIMS
[…]
MISA [13] - 76.5 76.6 0.447 0.563
TFN [60] 65.1 78.4 78.6 0.432 0.591
MulT [41] 64.8 78.6 79.7 0.453 0.564
MAG-BERT [36] - 74.4 71.8 0.492 0.399
DEVA [48] 65.4 79.6 80.3 0.424 0.583
KuDA[9] 66.5 80.7 80.7 0.408 0.613
HumanOmni [64] 72.9 85.1 85.0 0.327 0.749
U-ACS 73.1 85.8 85.1 0.354 0.691
IV-F Study on Computational Efficiency
In this subsection, we analyze the computational efficiency of HumanOmni. Table IV compares the computational costs, Acc2, and MAE of different approaches.
Compared with HumanOmni, our strategy achieves the optimal performance in the Acc2 metric on MOSI, MOSEI, and SIMS datasets, while attaining significant efficiency improvements. Specifically, on the MOSI dataset, the runtime of U-ACS is only 33% of the time HumanOmni takes, which significantly reduces the inference time while maintaining overall performance.
These results indicate that the proposed strategy, while ensuring that the model performance reaches the level of MLLM, has significantly shortened inference time and reduced model complexity, greatly improving operational efficiency and ultimately achieving a balance between performance and efficiency. We will further analyze the performance and adaptability of the proposed strategy across different MLLMs in Section IV-G.
TABLE IV: Computational efficiency comparison on HumanOmni. Time is measured in seconds per sample. Best results are in bold, second-best are underlined.
[…]
MOSI UBM 15.35 86.1 0.716
HumanOmni [64] 379.18 91.3 0.549
U-ACS 125.76 91.8 0.585
MOSEI UBM 35.02 86.3 0.527
HumanOmni [64] 9815.56 86.1 0.483
U-ACS 5561.18 86.4 0.506
SIMS UBM 18.76 80.5 0.409
HumanOmni [64] 470.61 85.1 0.327
U-ACS 260.58 85.8 0.354
IV-G Study and Analysis on Different MLLMs
To verify the generalization ability and robustness of the proposed U-ACS, we conduct experiments by integrating it with a different powerful MLLM: VideoLLaMA2. This subsection aims to demonstrate that the U-ACS strategy is not reliant on a specific model architecture but can effectively enhance the performance of different MLLM.
As shown in Table V, the U-ACS consistently improves Acc2 and F1-score for VideoLLaMA2 across the MOSI and SIMS datasets. For instance, when paired with VideoLLaMA2, U-ACS boosts the Acc2 score by 2.2 points on MOSI (from 88.3% to 90.5%) and a remarkable 4.6 points on SIMS (from 81.0% to 85.6%). A similar trend is observed with HumanOmni, where U-ACS improves the Acc2 by 0.5 points on MOSI and 0.7 points on SIMS. Notably, our method yields optimal Acc2 performance across all three datasets for VideoLLaMA2, which aligns with the findings for HumanOmni. While the standalone MLLMs occasionally perform better on other metrics such as MAE or Corr, the significant and consistent improvement in Acc2 and F1 highlights the effectiveness of our collaborative system. This outcome aligns with our design, which optimizes for binary polarity classification by using coarse-grained positive and negative categories for thresholding and uncertainty estimation. Moreover, in the UBM, we transform continuous sentiment labels into positive, negative, and neutral classes. These designs make the optimization of Acc2 (positive versus negative) more targeted, thereby highlighting significant advantages.
Meanwhile, the proposed strategy not only achieves optimal performance across both Acc2 and MAE on the SIMS dataset, but also reduces the computational time by nearly 50% on both the MOSI and SIMS datasets in Table VI. These results confirm that U-ACS is a model-agnostic framework that successfully leverages MLLMs to handle difficult samples, validating the broad applicability and effectiveness of our approach.
TABLE V: Generalization Evaluation on Different MLLMs. Best results are in bold, second-best are underlined.
[…]
MOSI VideoLLaMA2 [5] 49.8 88.3 88.3 0.575 0.875
U-ACS 48.9 90.5 90.5 0.625 0.847
SIMS VideoLLaMA2 [5] 74.0 81.0 81.7 0.400 0.718
U-ACS 74.4 85.6 85.6 0.392 0.685
TABLE VI: Computational efficiency comparison on VideoLLaMA2. Time is measured in seconds per sample. Best results are in bold, second-best are underlined.
[…]
MOSI UBM 15.35 86.1 0.716
VideoLLaMA2 [5] 1046.23 88.3 0.575
U-ACS 520.41 90.5 0.625
SIMS UBM 18.76 80.5 0.409
VideoLLaMA2 [5] 1351.04 81.0 0.400
U-ACS 710.71 85.6 0.392
IV-H Study on Hyperparameters
To analyze the impact of the threshold weighting hyperparameter $\lambda$ from Equation 19, we conducted a sensitivity analysis on MOSI and SIMS datasets. The hyperparameter $\lambda$ governs the uncertainty threshold, directly controlling how many samples are escalated to the MLLM. A lower $\lambda$ value results in a lower uncertainty threshold, causing more samples to be processed by the MLLM, while a higher $\lambda$ leads to a higher threshold, allowing the small model to handle more samples.
Figure 3 illustrates the trade-off between inference time and performance (Acc2 and MAE) as $\lambda$ varies from 0.1 to 1.0. The results reveal a clear trend: At lower values of the hyperparameter $\lambda$, more samples are escalated to the MLLM, meaning the MLLM needs significantly more time to infer these samples. Yet the MLLM is endowed with more robust inference capabilities, which in turn enables the system to deliver better performance on the two metrics (Acc2 and MAE). Thus, the primary factor affecting the system’s efficiency is the number of samples inferred by the MLLM. By adjusting the value of $\lambda$, we can find the optimal balance between inference time and model performance.
In Subsection IV-F, we elaborate on the efficiency of our system and demonstrate the trade-off between inference time and model performance. The hyperparameter $\lambda$ is therefore employed to identify an optimal value that maximizes the overall system performance. Consequently, when selecting hyperparameters, we need to ensure that the model achieves optimal performance across the target metrics. Therefore, for the main experiments in this study, we set the hyperparameter $\lambda$ to 0.1 for the MOSI dataset and to 0.2 for the SIMS dataset, respectively.
Refer to captionRefer to caption
(a) Acc2
Refer to captionRefer to caption
(b) MAE
Figure 3: Impact of the threshold weighting parameter $\lambda$ on inference time and two metrics (Acc2 and MAE). Top row: MOSI. Bottom row: SIMS.
IV-I Study on Uncertainty Distributions
To determine optimal uncertainty thresholds $\tau_{1}$ and $\tau_{2}$, We fit the uncertainty of correctly and incorrectly predicted samples to a Gaussian distribution. Figure 4 shows the uncertainty distributions for correct and incorrect predictions on MOSI testing set.
Although there is an overlap in the uncertainties of some samples, the uncertainty distributions of correctly predicted samples and incorrectly predicted samples on the MOSI dataset still show significant differences. The visualization indicates the effectiveness of the proposed uncertainty estimation method (especially for UBM whose uncertainty is hard to estimate). The medians of the two Gaussian distributions exhibit a significant difference, indicating that predicted samples and incorrectly predicted samples have distinct separability in terms of numerical features. Thus, we can optimize the threshold to find an optimal dividing point, which can maximizes the separation between predicted samples and incorrectly predicted samples, thereby achieving effective classification of predicted samples and incorrectly predicted samples.
(a) UBM
(b) MLLM
Figure 4: The distribution of uncertainties on MOSI.
IV-J Case Study
In this subsection, to illustrate the effectiveness of our strategy, we present a case study on a challenging sample from the SIMS dataset. Figure 5 demonstrates how our framework processes a complex sentiment analysis case where individual models struggle but collaboration system succeeds.
IV-K Ablation Studies
TABLE VII: Ablation experiments on MOSI and SIMS. Best results are in bold, second-best are underlined.
[…]
MOSI W/o Cross-Verification 50.4 90.4 0.631
W/o Uncertainty 51.7 89.6 0.600
Prediction-Truth Difference 46.7 87.0 0.649
Ensemble Variance 46.9 88.7 0.632
U-ACS 51.2 91.8 0.585
[…]
SIMS W/o Cross-Verification 72.9 85.1 0.358
W/o Uncertainty 70.7 83.6 0.357
Prediction-Truth Difference 70.2 81.2 0.407
Ensemble Variance 64.6 78.1 0.398
U-ACS 73.1 85.8 0.354
We conduct comprehensive ablation studies to validate each component of our framework. Table VII and Table VIII presents the results:
(3) Uncertainty Quantification Methods: Our entropy-based approach substantially outperforms alternative uncertainty estimation methods. Compared to prediction-truth difference, we achieve 4.8-point improvements in Acc2 on MOSI and 4.6-point improvements on SIMS, and achieves 0.014 decrease in MAE and a 2.0-point improve in Acc7 on MOSEI. Ensemble variance performs even worse, particularly on SIMS with a 7.7-point Acc2 gap and 9.2-point Acc5 gap. These results confirm that our single-pass entropy computation provides both efficiency and effectiveness for uncertainty estimation, which cleverly circumvents the issue of difficulty in computing uncertainty for regression task via transforming regression task into classification task.
TABLE VIII: Ablation experiments on MOSEI use HumanOmni as the MLLM. Best results are in bold, second-best are underlined.
[…]
MOSEI PTD 54.4 84.7 0.520
EV 55.2 86.4 0.516
CBE 55.4 86.4 0.506
IV-L Data Processing Volume in Different Components
Tables IX illustrate the actual volume of data processed across different components and datasets in our experiments. Owing to the varying levels of difficulty in sample processing across different datasets, the proportions of data routed to large and small models for processing also differ accordingly.
TABLE IX: Data Processing Volume (DPV) across different components of U-ACS. The MLLM utilized in this table is HumanOmni.
Dataset Test set size Method DPV
MOSI 686 UBM 447
HumanOmni 239
MOSEI 4659 UBM 1984
HumanOmni 2675
SIMS 457 UBM 278
HumanOmni 179
V Conclusion
We propose a novel method named U-ACS to combine UBM with MLLMs to address the imbalance between performance and efficiency in MSA. U-ACS explicitly computes the uncertainty of UBM and MLLM, designing an uncertainty-driven cascade mechanism to select samples yielding high predictive uncertainty for processing by the MLLM. We also propose advanced collaboration strategies to handle the hardest samples. U-ACS significantly improves the efficiency of MLLM while ensuring model performance across multiple datasets. Moreover, we show that the proposed strategy can generalize to different MLLMs.